OpenStudy (anonymous):

Solve the triangle: A = 51°, b = 14, c = 6

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@e.mccormick i just need a, B, and C

OpenStudy (loser66):



OpenStudy (loser66):

or else?

OpenStudy (amistre64):

law of cosines gives us \[a^2=b^2+c^2-2bc~cosA\] the rest is elementary

OpenStudy (anonymous):

Yeah thats it, I just need a C and B

OpenStudy (anonymous):

232 - 2(14)(6)cos51 = a^2 How do i do it?

OpenStudy (amistre64):

solving for the angle part: \[C=cos^{-1}\left(\frac{c^2-a^2-b^2}{-2ab}\right)\] \[B=cos^{-1}\left(\frac{b^2-a^2-c^2}{-2ac}\right)\]

OpenStudy (amistre64):

you generally do it with a calculator

OpenStudy (anonymous):

but how do i do this? 232 - 2(14)(6)cos51 = a^2

OpenStudy (anonymous):

Thanks! Just that one value gave me the other ones.

OpenStudy (amistre64):

:)